OpenStudy (lilshane):

Which three ordered pairs complete the table to give solutions of the equation y = x – 5? A. (–3, –2), (0, –5), and (3, 2) B. (–3, –2), (0, –5), and (3, –2) C. (–3, –8), (0, –5), and (3, 2) D. (–3, –8), (0, –5), and (3, –2)

OpenStudy (lilshane):

@texaschic101

OpenStudy (lilshane):

@iGreen

OpenStudy (lilshane):

@nincompoop

OpenStudy (lilshane):

@ortiz09

OpenStudy (lilshane):

@Compassionate @TheSmartOne @aaronq

OpenStudy (lilshane):

@confluxepic @paki

OpenStudy (anonymous):

hold on im checking

OpenStudy (lilshane):

@e.mccormick

OpenStudy (anonymous):

D

OpenStudy (lilshane):

back up the answer please

OpenStudy (anonymous):

-8=-3-5 -8=(3)+(-5) -8=-8 which would then give you the total of 5

OpenStudy (anonymous):

(-3,-8)

OpenStudy (anonymous):

so it would be D